OpenStudy (anonymous):

an executive committee consists of 12 members: 5 men and 7 women. 3 members are selected at random to attend a meeting in Hawaii. The names are drawn from a hat. What is the probability that all 3 selected are men?

OpenStudy (anonymous):

\[\frac{ 5 }{ 7} x \frac{ 4 }{ 6} x \frac{ 3 }{ 5 }\]

OpenStudy (anonymous):

Fine the answer to this.

OpenStudy (anonymous):

Find*

OpenStudy (anonymous):

DANGGG!! Sorry I misread, 7 is the number of woemn!

OpenStudy (anonymous):

\[\frac{ 5 }{ 12 } x \frac{ 4 }{ 11 } x \frac{ 3 }{ 10 }\]

OpenStudy (anonymous):

Sorry for the wrong answer.